# Information theory for data-driven model reduction in physics and biology

### v2

@@ II Information decay and the spectrum of the transfer operator
Figure 2: Information loss of a Brownian particle in a double well potential. (a) Fixed point (FP) diagram of the dynamics given by Eq. 7 for zero noise. There is a bifurcation at $\mu=0$ where the stable FP at $x=0$ becomes unstable and two new stable FPs appear at ±$\pm\sqrt{\mu}$. Insets show the evolution of the corresponding potential $V(x)$, with the emergence of a double-well structure for $\mu>0$. (b) Dynamics of the system Eq. 7 for varying values of $\mu$ corresponding to the potential insets in (a), with uniformly-distributed initial conditions. (c) Loss of information between the initial condition and the future state. Inset shows scaling given by the first eigenvalue of the transfer operator. (d) Mutual information between the present and future state for varying time delay $\Delta t$ and bifurcation parameter $\mu$. (e) Spectrum of the transfer operator $U$, showing a pile-up of eigenvalues for $\mu\gtrsim 0$. These are related to the eigenvalues of its infinitesimal generator by $\Lambda_{i}=\text{e}^{\lambda_{i}\Delta t}$. (f) Maximal mutual information which can be encoded into a discrete variable of $N_{H}$ values, for a fixed time delay $\Delta t=1.0$. Black dashed line shows $I(X_{t},X_{t+\Delta t})$ for reference. Information is provided in units of bits.
@@
To develop intuition for information in a dynamical system, we turn to the simple example of a Brownian particle trapped in a confining double-well potential. This might represent, for example, a molecule with a single degree of freedom that transitions between two metastable configurations [42]. In the overdamped limit the state of the particle is completely determined by its position $X_{t}\in\mathbb{R}$, with dynamics given by the Langevin equation
@@
The deterministic dynamical system undergoes a bifurcation at $\mu=0$ (Fig. 2a). Sample trajectories, with noise, for a uniform initial distribution of particles are shown in Fig. 2b. For negative $\mu$, the trajectories all converge to a fixed point at $x=0$, while for $\mu>0$ they fluctuate around the fixed points at $x=\pm\sqrt{\mu}$.
To quantify the amount of information about the future state $X_{t+\Delta t}$ contained in the initial state $X_{t}$ we compute their mutual information (Fig. 2c; see SI for details). The dynamics of $X_{t}$ are Markovian, so that for any sequence of times $t_{0}<t_{1}<t_{2}$, $p(X_{t_{2}}|X_{t_{1}},X_{t_{0}})=p(X_{t_{2}}|X_{t_{1}})$. From the data processing inequality, one has [43]
@@
where expectations are taken over the steady state distribution (see SI). Asymptotically, the information decay is set by the value of $\lambda_{1}$, the rate of decay of the slowest-varying function $\phi_{1}(x)$ under the dynamics of $U$. In the limit of infinite time, for any value of $\mu$ even weak noise will cause the mutual information to become zero as there is a non-zero probability of hopping between the wells, though this rate of hopping is exponentially small [33].
@@
The loss of information in time depends on the bifurcation parameter $\mu$ as summarized in Fig. 2d. Note the peak in $I(X_{t},X_{t+\Delta t})$ for small, positive $\mu$. This corresponds to dynamics where observation of $X_{t}$ strongly informs the future state; recall that the mutual information is maximized when the conditional entropy $S(X_{t+\Delta t}|X_{t})\approx 0$ (see Methods). In contrast, for large positive or negative $\mu$, $X_{t}$ is not as informative of $X_{t+\Delta t}$ even for small times: the initial state is quickly forgotten as the particle approaches the bottom of the single (for $\mu<0$) or double (for $\mu>0$) well.
@@
This phenomenon is reminiscent of critical slowing down, which occurs in the noise-free system as $\mu$ passes through the bifurcation at $\mu=0$. For the deterministic dynamics, the slowing down is reflected in the spectrum as a “pile up” of eigenvalues to form a continuous spectrum [33]. In the presence of noise, although the continuous spectrum becomes discrete [33, 34] there is still a pile-up of eigenvalues characterized by several eigenvalues becoming close to 1 (Fig. 2e). This pile-up gives rise to the information peak seen in Fig. 2d. The peak is not solely due to the closing spectral gap $\lambda_{1}-\lambda_{2}$, but is also impacted by the subdominant eigenvalues which accumulate at $\mu\approx 0.2$ (SI Fig. S4).
III Knowing when to stop
@@
For discrete encoding variables $h$, the information bottleneck partitions state space and reduces the dynamics on $x$ to a discrete dynamics on $h$. Such reductions of complex systems to symbolic sequences via partitioning of state space has attracted attention for more than half a century in both theoretical and data-driven contexts[44, 45, 46, 47, 48, 49, 50]. Several works have approached this partition problem from a dynamical systems perspective, linking “optimal” partitions to eigenfunctions of the (adjoint) transfer operator [25, 51]. In this setting, a central question is “when to stop” [25, 26, 48, 49]: how many states does $h$ need in order to capture statistical properties of the original dynamics?
@@
We consider this question by finding the optimal IB encoder in the limit of low compression, $\beta\gg 1$, but fixed encoding capacity $N_{H}$ (where $H_{t}\in\{0,...,N_{H}-1\}$), i.e. the encoder is only restricted by the number of symbols it can use. An analogous setup was used in the context of renormalization group (RG) transformations in [52, 53, 54], which results in effective model reduction due to the “sloppiness”, or irrelevance, of certain system variables [55, 56]. In this regime, the encoder learned by IB is deterministic; we are learning an optimal hard partition of state space. This can be seen by noting that $I(H_{t};X_{t+\Delta t})=S(H_{t})-S(H_{t}|X_{t+\Delta t})$ is maximized when the latter term is zero, which happens when $x_{t}$ unambiguously determines $h_{t}$, i.e. when $p(h_{t}|x_{t})\in\{0,1\}$ for all $x_{t}$. The details of how the encoder is computed are discussed in the next section.
Fig. 2f shows that the number of states necessary to describe the system depends strongly on the value of $\mu$. For $|\mu|\gg 0$, a two-state discrete variable $h_{t}\in\{0,1\}$ suffices to describe the system’s future. Increasing the number of reduced variables $N_{H}$ does not allow more information to be captured. Near the information peak at $\mu\approx 0.2$ this changes: predicting the future state of the system requires a more complex hidden variable of up to $N_{H}\approx 10$ values. Above, we saw that this peak arises due to the pile-up of eigenvalues at $\mu\approx 0.2$. The content of the transfer operator spectrum is thus directly reflected in the number of encoding variables needed to capture the system’s statistics.
Noise can have a similarly dramatic impact on the reduced model complexity. Indeed, noise in some form, either inherent to the dynamics or due to measurement error, is necessary for a model to be reducible. In purely deterministic systems where the future state is a bijective function of the present state, information does not decay and complete knowledge of the state is required to predict the future.
@@
Consider a fluctuating Brownian particle as above, where now each of the wells is split into two smaller wells, giving a total of four potential minima (Fig. 3a). As the system is in steady state, the standard deviation of the fluctuations $\sigma$ corresponds to an energy scale $E_{\sigma}=\sigma^{2}=2k_{\text{B}}T$. For small $E_{\sigma}$, the system rarely transitions between the four potential minima. In this case, knowledge of the initial minimum is very informative of the future state of the particle. In contrast, for large fluctuations the particle can spontaneously jump between shallow minima in each large well, so that the system immediately forgets about the precise potential minimum it was in. Information about the shallow minima has been “washed out”, and only the information about the larger double-well structure remains.
@@
To see this reflected in the information, we again consider an encoding of the initial state into a discrete variable $H_{t}\in\{0,...,N_{H}-1\}$. In both the small and large noise scenarios, a variable with $N_{H}=2$ encodes approximately one bit of information (Fig. 3b), corresponding to an $H_{t}$ which distinguishes the two large wells for $x\lessgtr 0$. For large noise this is essentially all the information that can be learned; increasing the capacity of the encoding variable beyond this provides only marginally more information about the future state (Fig. 3c). In the small noise case, the information between the encoding and the future state continues to increase to approximately two bits at $N_{H}=4$, after which it plateaus. The encoding has learned to distinguish each of the four potential wells. These observations are reflected in the transfer operator spectrum shown in Fig. 3d. For small noise, the eigenvalue $\lambda=1$ is four-fold degenerate which indicates the existence of four regions that can evolve independently under $U$, giving rise to four steady state distributions satisfying $U\rho=\rho$. These regions correspond to the potential minima. Hops between the separate minima are exceedingly rare, so that the dynamics essentially take place in the four minima independently. With larger $\sigma$ the degeneracy is lifted, resulting in one dominant subleading eigenvalue followed by a gap. The corresponding eigenfunction is one which is positive (negative) on the right (left) side of the large potential barrier at $x=0$: the only relevant piece of information is which of the large wells the initial condition is contained in, and all other information is lost exponentially quickly.
IV Transfer operator eigenfunctions are most informative features
Until now we have concerned ourselves with encodings whose capacity is limited only by the number of variables, rather than by the compression imposed by a small value of $\beta$. In the regime of small $\beta$, or high compression, features of the state $x_{t}$ are forced to compete to make it through the bottleneck $h_{t}$. By studying the behavior of the encoder in this regime, in particular its dependence on $x_{t}$, we may identify the most relevant features of the state variable and show that they coincide with left eigenfunctions of the transfer operator.
We return to the simple example of a particle in a double well with dynamics given by (7) which we map to a discrete variable $H_{t}\in\{0,...,N_{H}-1\}$. In this system the IB loss function (1) can be optimized directly, as shown in Ref. [30], using an iterative scheme known as the Blahut-Arimoto algorithm [43] (see SI).
To focus on the properties of encodings for varying degrees of compression $\beta$, we consider a fixed set of dynamical parameters $\mu$ and $\sigma$. Increasing $\beta$ reduces the amount of compression, i.e. “widens” the bottleneck, allowing more information to pass into the encoder. This leads to a series of IB transitions which are sketched in Fig. 1b and shown quantitatively in Fig. 4a. The form of the optimal encoder changes qualitatively at these transitions. Before $\beta_{1}$, the optimal encoder has no dependence on $x$ so that $p(H_{t}=h_{i}|x_{t})=\text{const}$ for all $h_{i}$. After the first transition, the encoder begins to associate regions of $x$ to particular values of $h$. We are interested in the form of the encoder at $\beta\gtrsim\beta_{1}$, just above the first IB transition, as this reflects the most informative features of the full state variable $x$ (Fig. 4a). The dependence of $p(h_{t}|x_{t})$ on $x$ can be explained by a stability analysis of the IB Lagrangian (see Methods and SI). Stability is governed by the eigenvalues $\eta_{i}$ of the Hessian of the IB Lagrangian with respect to the parameters $f_{n}(h_{t})$ in Eq. 5. These parameters tell us how much the encoder “weights” each transfer operator eigenfunction; $f_{n}(h_{t})=0$ (for $n>0$) corresponds to the uniform, or trivial encoder $p(h_{t}|x_{t})=p(h_{t})$.
Figure 4: IB learns eigenfunctions of the adjoint transfer operator. (a) When the relative weight $\beta$ between both constraints in Eq. 1 is changed, more and more information can go through the encoder. This occurs in steps, where the spectral content of the transfer operator is included starting from eigenvalues with largest magnitude (i.e., the slowest ones). (b) Transitions are characterized by the appearance of negative eigenvalues in the spectrum of the Hessian of the IB loss function. Here we consider the Hessian evaluated at the uniform encoder $p(h|x)=N_{H}^{-1}$. The IB transitions $\beta_{1}\approx 1.1$ and $\beta_{2}\approx 1.8$ correspond to the appearence of negative eigenvalues of the Hessian. (c) The unstable directions are dominated by single components (note the color scale is logarithmic). (d) At the first transition, the logarithm of the encoder is given by the eigenfunction $\phi_{1}(x)$, up to rescaling ($y$-axis is shown in arbitrary units). (e) Likewise, at the second transition the encoder is given by $\phi_{2}(x)$.
For small $\beta$ all eigenvalues $\eta_{i}$ are positive, indicating that the uniform encoder is a stable minimum of the IB Lagrangian. In Fig. 4b we show the smallest two eigenvalues of the IB Hessian when evaluated at the uniform encoder. At the first transition one eigenvalue becomes negative, so that the uniform encoder is unstable. The eigenvector corresponding to the unstable eigenvalue $\eta_{1}$ indicates how the weights $f_{n}(h_{t})$ should be adjusted to lower the value of the IB Lagrangian. Our numerics confirm that these weights are dominated by $f_{1}$ as expected from our analytical result (Fig. 4c, top). By taking the logarithm of the encoder after the transition, we can independently confirm that the encoder depends only on $\phi_{1}(x)$ (Fig. 4d).
Our stability analysis predicts that a second mode becomes unstable at the second IB transition $\beta\approx\beta_{2}$ (Fig. 4b). Here we see that this unstable mode selects $f_{2}$, and that the encoder correspondingly gains dependence on $\phi_{2}(x)$ (Fig. 4d). Note that in general, $\eta_{2}$ must not necessarily become negative precisely at $\beta_{2}$ because the stability analysis is performed at the uniform encoder while the true optimal encoder has already deviated from uniformity. In the SI, we perform the same analysis for a triple-well potential where this difference is more apparent.
@@ V Data-driven discovery of slow variables
IB finds transfer operator eigenfunctions by optimizing an information theoretic-objective that makes no reference to physics or dynamics. This suggests it may be used for the discovery of slow variables in situations where one lacks physical intuition. The utility of exact IB for this purpose is limited because it requires knowledge of the exact conditional distribution $p(x_{t+\Delta t}|x_{t})$ which is difficult to estimate in many real-world scenarios. Fortunately however, the IB optimization problem can be replaced by an approximate variational objective introduced in Ref. [41] that can be solved with neural networks. We refer to this as variational IB. In the remainder of this paper, we show how to implement these networks for the discovery of slow variables directly from data.
@@
Figure 5: Variational IB for high-dimensional simulated fluid flow. (a) A fluid flows into the system with uniform velocity $v_{0}$ in the $x$-direction and passes a disk-shaped obstacle, which perturbs the fluid and causes vortex shedding behind the object in a so-called von Kármán street. The state of the system is given by a spatially varying two-component vector field $\mathbf{v}(\mathbf{x})$. (b) The dynamics in latent space (blue) are very regular, traversing a nearly circular trajectory (see Supplementary Movie 1). For comparison we show the evolution of the mode amplitudes obtained by projecting the velocity field onto the first DMD mode (green). (c) Time evolution of one component of the latent variable ($h_{1}$, blue) as well as the DMD mode amplitude (green). (d) Comparison of the first Koopman mode obtained from DMD ($\mathbf{m}^{(1)}$) and from VIB ($\mathbf{m}^{(\text{IB})}$). Koopman modes from VIB are computed as gradients of the latent encoding variables as described in the main text. Red corresponds to positive values and blue to negative; the magnitudes of the modes are not directly comparable.
@@
First we show numerically that the results of the previous sections remain valid for high-dimensional systems by considering a simulated data of fluid flow past a disk [57]. The state of the system is given by a two-dimensional velocity field $\mathbf{v}(\mathbf{x})\in\mathbb{R}^{2\times N_{\text{pixels}}}$, where $N_{\text{pixels}}\sim\mathcal{O}(10^{5})$ (Fig. 5a). Fluid flows in from the left boundary with a constant velocity $v_{0}\hat{e}_{x}$ past a disk of unit diameter. At Reynolds number $\text{Re}\gtrsim 150$, the fluid undergoes periodic vortex shedding behind the disk, forming what is known as a von Kármán street.
$\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle,$ (9)
@@
where $\mathbf{m}^{(n)}$ is the $n$-th mode (often referred to as a Koopman mode [19]) and angled brackets denote integration over space. The true eigenfunction and corresponding modes can be computed via dynamic mode decomposition (DMD) [11, 12], as described in the SI. The eigenfunctions for this system are in general complex, and come in conjugate pairs: $\phi_{2}(x)=\phi^{*}_{1}(x)$. In this situation any linear combination of $\phi_{1}$ and $\phi_{2}$ will decay at the same rate, and hence we expect to learn some arbitrary combination of the two dominant eigenfunctions, or equivalently a combination of the real and imaginary parts of $\phi_{1}$. We therefore take a two-dimensional encoding variable $[h_{0},h_{1}]$, so that it can represent the full complex eigenfunction rather than only the real or imaginary part.
Our learned latent variables are oscillatory with the correct frequency as shown in Fig. 5b-c. A more stringent test is whether we are also learning the correct mode $\mathbf{m}^{(1)}$. From the true eigenfunctions, the modes can be extracted by computing the gradient
$\frac{\partial\phi_{n}}{\partial v_{j}}=m_{j}^{(n)}.$ (10)
$\frac{\partial h}{\partial v_{j}}=m^{(\text{IB})}_{j}+g_{\text{res},j}(\mathbf{v}(\mathbf{x})),$ (11)
where we have separated the part of the gradient which is independent of $\mathbf{v}$ from a residual part which is dependent on $\mathbf{v}$. If $h$ corresponds to the true eigenfunction, we expect that $\mathbf{m}^{(\text{IB})}$ is approximately equal to the Koopman mode $\mathbf{m}^{(1)}$, and that $\mathbf{g}_{\text{res}}$ is small. We indeed find this to be the case; Fig. 5d shows these gradients averaged over several instantiations of the neural network, which corresponds strongly to the true mode. Details concerning both the averaging procedure and the residuals $\mathbf{g}_{\text{res}}$ can be found in the SI. This shows that variational IB not only recovers the essential oscillatory nature of the dynamics, but does so by learning the correct slowly varying functions of the state variable given by the adjoint transfer operator eigenfunctions.
VI Relevant variable identification in laboratory-generated and atmospheric flows
The scenario above is characterized by high-dimensional data and few samples; training was performed with only $\sim 400$ samples. We now demonstrate that our framework continues to hold approximately and yield interpretable latent spaces even for real-world fluid flow datasets scraped directly from videos on Youtube [58, 59] (Supplementary Movie 1).

### v3

@@ III Information decay and the spectrum of the transfer operator
Figure 2: Information loss of a Brownian particle in a double well potential. (a) Fixed point (FP) diagram of the dynamics given by Eq. 9 for zero noise. There is a bifurcation at $\mu=0$ where the stable FP at $x=0$ becomes unstable and two new stable FPs appear at $\pm\sqrt{\mu}$. (a,i) The corresponding potential $V(x)$ for varying $\mu$, with the emergence of a double-well structure for $\mu>0$. (a,ii) Dynamics of the system Eq. 9 for varying values of $\mu$ corresponding to the potentials above, with uniformly-distributed initial conditions. (a,iii) Loss of information between the initial condition and the future state. Inset shows scaling given by the first eigenvalue of the transfer operator. (b) Spectrum of the transfer operator $U$, showing a pile-up of eigenvalues for $\mu\gtrsim 0$. These are related to the eigenvalues of its infinitesimal generator by $\Lambda_{i}=\text{e}^{\lambda_{i}\Delta t}$. (c) Mutual information between the present and future state for varying time delay $\Delta t$ and bifurcation parameter $\mu$.
@@
Figure 4: Non-Markovianity for a Brownian particle in a double well potential. (a) Conditional mutual information $I(H_{t-\Delta t},H_{t+\Delta t}|H_{t})$ as a function of $N_{H}$. (b) Two possible scenarios (left and right) with the same value of $h_{t}$, but different $h_{t-1}$. Here, the circles represent the value of the full state $x_{t}$ while the colored regions represent the value of the reduced state $h_{t}\in\{1,2\}$ which are mapped to for each $x_{t}$. (c) Illustration of the transition probability which takes into account the previous state $h_{t-1}$ (left), compared to the Markovian transition probability obtained by marginalizing out $h_{t-1}$ (right). In a Markovian system, these transition probabilities would coincide. (d-e) Same as in (h), but for $N_{H}=4$ states.
To develop intuition for information in a dynamical system, we turn to the simple example of a Brownian particle in a double-well potential. This may represent, for example, a molecule with a single degree of freedom that transitions between two metastable configurations [57]. In the overdamped limit the state of the particle is completely determined by its position $X_{t}\in\mathbb{R}$, with dynamics given by the Langevin equation
@@
The deterministic dynamical system undergoes a bifurcation at $\mu=0$ (Fig. 2a). To quantify the amount of information about the future state $X_{t+\Delta t}$ contained in the initial state $X_{t}$ we compute their mutual information (see SI for details). The dynamics of $X_{t}$ are Markovian, so that for any sequence of times $t_{0}<t_{1}<t_{2}$, $p(X_{t_{2}}|X_{t_{1}},X_{t_{0}})=p(X_{t_{2}}|X_{t_{1}})$. From the data processing inequality, one has [58]
@@
where expectations are taken over the steady state distribution (see SI Section 3). Asymptotically, the information decay is set by the value of $\lambda_{1}$, the rate of decay of the slowest-varying function $\phi_{1}(x)$ under the dynamics of $U$. In the limit of infinite time, for any value of $\mu$ even weak noise will cause the mutual information to become zero as there is a non-zero (perhaps exponentially small) probability of hopping between the wells [49].
@@
The loss of information in time depends on the bifurcation parameter $\mu$ as summarized in Fig. 2c. Note the peak in $I(X_{t},X_{t+\Delta t})$ for small, positive $\mu$. This corresponds to dynamics where observation of $X_{t}$ strongly informs the future state; recall that the mutual information is maximized when the conditional entropy $\mathcal{H}(X_{t+\Delta t}|X_{t})\approx 0$ (see SI Section 1). In contrast, for large positive or negative $\mu$, $X_{t}$ is not as informative of $X_{t+\Delta t}$ even for small times: the initial state is quickly forgotten as the particle approaches the bottom of the single (for $\mu<0$) or double (for $\mu>0$) well.
@@
This phenomenon is reminiscent of critical slowing down, which occurs in the noise-free system as $\mu$ passes through the bifurcation at $\mu=0$. For the deterministic dynamics, the slowing down is reflected in the spectrum as a “pile up” of eigenvalues to form a continuous spectrum [49]. In the presence of noise, although the continuous spectrum becomes discrete [49, 50] there is still a pile-up of eigenvalues characterized by several eigenvalues becoming close to 1 (Fig. 2b). This pile-up gives rise to the information peak seen in Fig. 2c. The peak is not solely due to the closing spectral gap $\lambda_{1}-\lambda_{2}$, but is also impacted by the subdominant eigenvalues which accumulate at $\mu\approx 0.2$ (SI Fig. S2).
IV Knowing when to stop
@@
For discrete encoding variables $h$, the information bottleneck partitions state space and reduces the dynamics on $x$ to a discrete dynamics on $h$. Such reductions of complex systems to symbolic sequences via partitioning of state space has attracted attention for more than half a century in both theoretical and data-driven contexts [59, 60, 61, 62, 4, 63, 64]. Several works have approached this partition problem from a dynamical systems perspective, linking optimal partitions to eigenfunctions of the (adjoint) transfer operator [33, 65]. In this setting, a central question is “when to stop” [33, 34, 4, 63]: how many states does $h$ need in order to capture statistical properties of the original dynamics?
@@
We consider this question by finding the optimal IB encoder in the limit of low compression, $\beta\gg 1$, but fixed encoding capacity $N_{H}$ (where $H_{t}\in\{0,...,N_{H}-1\}$), i.e. the encoder is only restricted by the number of symbols it can use. An analogous setup was used in the context of renormalization group (RG) transformations in [66, 67, 68], which results in effective model reduction due to the “sloppiness”, or irrelevance, of certain system variables [69, 70]. In this regime, the encoder learned by IB is deterministic; we are learning an optimal hard partition of state space. This can be seen by noting that $I(H_{t};X_{t+\Delta t})=\mathcal{H}(H_{t})-\mathcal{H}(H_{t}|X_{t+\Delta t})$ is maximized when the latter term is zero, which happens when $x_{t}$ unambiguously determines $h_{t}$, i.e. when $p(h_{t}|x_{t})\in\{0,1\}$ for all $x_{t}$. The details of how the encoder is computed are discussed in the next section.
@@
Consider a fluctuating Brownian particle as in the double well above, where now each of the wells is split into two smaller wells, giving a total of four potential minima (Fig. 3a). As the system is in steady state, the variance of the fluctuations defines an energy scale $E_{\sigma}=\sigma^{2}=2k_{\text{B}}T$. For small $E_{\sigma}$, the system rarely transitions between the four potential minima. In this case, knowledge of the initial minimum is very informative of the future state of the particle. In contrast, for large fluctuations the particle can spontaneously jump between shallow minima in each large well, so that the system immediately forgets about the precise potential minimum it was in. Information about the shallow minima has been “washed out”, and only the information about the larger double-well structure remains.
@@
To see this reflected in the information, we again consider an encoding of the initial state into a discrete variable $H_{t}\in\{0,...,N_{H}-1\}$. In both the small and large noise scenarios, a variable with $N_{H}=2$ encodes approximately one bit of information (Fig. 3b), corresponding to an $H_{t}$ which distinguishes the two large wells for $x\lessgtr 0$. For large noise this is essentially all the information that can be learned; increasing the capacity of the encoding variable beyond this provides only marginally more information about the future state (Fig. 3c). In the small noise case, the information between the encoding and the future state continues to increase to approximately two bits at $N_{H}=4$, after which it plateaus. The encoding has learned to distinguish each of the four potential wells. These observations are reflected in the transfer operator spectrum shown in Fig. 3d. For small noise, the eigenvalue $\lambda=1$ is nearly four-fold degenerate, indicating the existence of four regions that can evolve independently under $U$, giving rise to four steady state distributions satisfying $U\rho=\rho$. These regions correspond to the potential minima. Hops between the separate minima are exceedingly rare, so that the dynamics essentially take place in the four minima independently. With larger $\sigma$ the degeneracy is lifted, resulting in one dominant subleading eigenvalue followed by a gap. The corresponding eigenfunction is one which is positive (negative) on the right (left) side of the large potential barrier at $x=0$: the only relevant piece of information is which of the large wells the initial condition is contained in, and all other information is lost exponentially quickly.
In many cases, our goal is not only to find a reduced representation $H_{t}$ that is predictive of the future state, but one which may be evolved forward for long times as a surrogate for the full system. This is a practical but much more stringent requirement of our reduced state, and it may result in a different answer to the question of “when to stop.”
In general the dynamics of the reduced variable will be non-Markovian, meaning that its evolution will be determined not only by its current state, but also its past. With a fixed encoder $p(h_{t}|x_{t})$, transition probabilities $p(h_{t+1}|h_{t})$ are induced by decoding $h_{t}$ to $x_{t}$, evolving $x_{t}$ forward in time to $x_{t+1}$, and then encoding $x_{t+1}$ to $h_{t+1}$ (here we assume discrete time dynamics with $\Delta t=1$ for simplicity). Information theory provides a natural quantification of the (non-)Markovianity of the dynamics as measured by the conditional mutual information $I(H_{t-1},H_{t+1}|H_{t})$ (see SI Section 3 and alternative approaches in Refs. [71, 72, 63]). This quantity tells us how much additional information the past state $H_{t-1}$ contains about the future $H_{t+1}$, on top of what is contained in the present state $H_{t}$.
To illustrate this, we again consider a particle in a double well, which we compress into a discrete variable $H_{t}\in\{0,...,N_{H}-1\}$. Because the full system is Markovian by design, we expect that for large $N_{H}$ the conditional mutual information $I(H_{t-1},H_{t+1}|H_{t})$ should go to zero as $H_{t}$ begins to more closely approximate the full state $X_{t}$ (Fig. 4a). Interestingly, the approach to Markovianity is highly non-monotonic, and the system is more non-Markovian with $N_{H}=4$ than with $N_{H}=2$.
The role of memory is illustrated in the two cases shown in Fig. 4b. In both cases, the current state is in the right well, so that $h_{t}=1$. However, more precise information about the state (and hence its future) is revealed by knowing the system’s past. If $h_{t-1}=0$ (so the state was in the left well), it is likely that the full state is near the boundary between wells, so that $h_{t+1}=0$ is more likely than if $h_{t-1}=1$, which would suggest that the state is deep in the right well. This behavior can be quantified by computing the transfer probabilities $p(h_{t+1}|h_{t},h_{t-1})$ (Fig. 4c). These differ significantly from $p(h_{t+1}|h_{t})$, which would not be the case for Markovian dynamics. From these, we see that the primary contribution to the non-Markovian dynamics occurs when $(h_{t-1},h_{t})=(0,1)$ or $(1,0)$, which can only happen when the state is very near to the boundary between the wells. As $N_{H}$ increases, the likelihood that the state is near the boundary of two regions with different $h_{t}$ labels will increase, which will increase the non-Markovianity of the system (Fig. 4d-e). However, at the same time the different $h_{t}$ regions will shrink in size so that $h_{t-1}$ will no longer provide any meaningful information about the system’s location $x_{t}$. The trade-off between these two effects leads to the observed non-monotonic behavior of $I(H_{t-1},H_{t+1}|H_{t})$.
@@ V Data-driven discovery of slow variables
IB finds a reduced state variable by optimizing an information theoretic-objective that makes no reference to physics or dynamics. This suggests it may be used for the discovery of slow variables in situations where one lacks physical intuition. In the regime of small $\beta$, or high compression, features of the state $x_{t}$ are forced to compete to make it through the bottleneck $h_{t}$, which allows us to extract relevant features of the system.
These variables coincide with left eigenfunctions of the transfer operator, as discussed in Section II. We verify this theoretical result numerically in cases where the IB objective (3) can be solved exactly by using an iterative scheme known as the Blahut-Arimoto algorithm [38, 58] (see SI Section 3). By taking the logarithm of the encoder as it begins to deviate from the trivial uniform encoder above $\beta_{1}$ we can confirm that it depends on $x$ only through $\phi_{1}(x)$ (SI Fig. S3). This can be independently verified by studying the stability of the trivial encoder with respect to perturbations in the parameters $f_{n}(h_{t})$ in Eq. 7, where we see that the encoder becomes unstable to perturbations in $f_{1}$, as predicted by our theoretical results (SI Section 4).
@@
The utility of exact IB for variable discovery is limited because it requires knowledge of the exact conditional distribution $p(x_{t+\Delta t}|x_{t})$ which is difficult to estimate in many real-world scenarios. Fortunately however, the IB optimization problem can be replaced by an approximate variational objective introduced in Ref. [41] that can be solved with neural networks and which we refer to as variational IB. This is achieved by introducing a tractable Ansatz for the form of $p(h_{t}|x_{t})$; we show in SI Section 4 that this does not change the encoder’s dependence on transfer operator eigenfunctions.
@@
Figure 5: Variational IB for high-dimensional simulated fluid flow. (a) Fluid flow past a disk with uniform in-flow velocity $v_{0}$ exhibits vortex shedding behind the object in a so-called von Kármán street. The state of the system is given by a spatially varying two-component vector field $\mathbf{v}(\mathbf{x})$. (b) Dynamics in latent space (blue) traverse a nearly circular trajectory (see Supplementary Movie 1). For comparison we show the evolution of the mode amplitudes obtained by projecting the velocity field onto the first DMD mode (green). (c) Comparison of the first Koopman mode obtained from DMD ($\mathbf{m}^{(1)}$) and from VIB ($\mathbf{m}^{(\text{IB})}$). Koopman modes from VIB are computed as gradients of the latent encoding variables as described in the main text. Red corresponds to positive values and blue to negative; the magnitudes of the modes are not directly comparable. (d) Intensity field $I(x)$ from a video of water flowing past an obstacle; video from Ref. [73]. (e) Trajectory of the system in latent space. (f) Gradients of the latent encodings with respect to the input field $I(x)$. (g-i) Same as (d-f), but using satellite images of clouds near the island of Guadalupe as our state $I(x)$.
@@
Our core result linking the behavior of the learned encoder with transfer operator eigenfunctions remains valid even for variational IB applied to high-dimensional systems, which we show by considering a simulated data of fluid flow past a disk [74]. The state of the system is given by a two-dimensional velocity field $\mathbf{v}(\mathbf{x})\in\mathbb{R}^{2\times N_{\text{pixels}}}$, where $N_{\text{pixels}}\sim\mathcal{O}(10^{5})$ (Fig. 5a). Fluid flows in from the left boundary with a constant velocity $v_{0}\hat{e}_{x}$ past a disk of unit diameter. At Reynolds number $\text{Re}\gtrsim 150$, the fluid undergoes periodic vortex shedding behind the disk, forming what is known as a von Kármán street.
@@
The eigenfunctions for this system are in general complex, and come in conjugate pairs: $\phi_{2}(x)=\phi^{*}_{1}(x)$. In this situation any linear combination of $\phi_{1}$ and $\phi_{2}$ will decay at the same rate, and hence we expect to learn some arbitrary combination of the two dominant eigenfunctions, or equivalently a combination of the real and imaginary parts of $\phi_{1}$. We therefore take a two-dimensional encoding variable $h_{0},h_{1}]$, so that it can represent the full complex eigenfunction.
Our learned latent variables are oscillatory with the correct frequency as shown in Fig. 5b. However, this alone does not indicate whether we are learning the correct eigenfunctions. Because the system is well approximated by linear dynamics, eigenfunctions of the adjoint transfer operator are linear functions of the state variable, and may hence be represented as $\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle$, where angled brackets denote an average over space and $\mathbf{m}^{(n)}(\mathbf{x})$ are so-called Koopman modes. The true $\mathbf{m}^{(n)}$ (and hence the eigenfunctions) can be thus computed via dynamic mode decomposition (DMD) [15, 16] as described in the SI. For the functions $h[\mathbf{v}]$ learned using variational IB, we can infer the corresponding $\mathbf{m}^{(\text{IB})}$ using gradients of the neural network with respect to the input field. Fig. 5c shows these inferred modes $\mathbf{m}^{(\text{IB})}$ compared to the true mode $\mathbf{m}^{(1)}$ (see SI Section 5 for details). This shows that variational IB not only recovers the essential oscillatory nature of the dynamics, but does so by learning the correct slowly varying functions of the state variable given by the adjoint transfer operator eigenfunctions.